OpenStudy (anonymous):

Can any one tell cube root of 456533 along the solution please

OpenStudy (neer2890):



OpenStudy (neer2890):

we take pair of 3 as one. so 7*7*7 is one pair. similarly, 11*11*11 so\[\sqrt[3]{456533}=7*11=77\]